# A Comparison of DeepSeek and Other LLMs

### v2

@@ 1 Introduction
In the past two weeks, DeepSeek (DS), a recent large language model (LLM) (DeepSeek-AI, 2024, 2025), has shaken up the AI industry. Since its latest version was released on January 20, 2025, DS has made the headlines of news and social media, shot to the top of Apple Store’s downloads, stunning investors and sinking some tech stocks including Nvidia.
@@
  • In Evstafev (2024), the author showed that over $30$ challenging mathematical problems derived from the MATH dataset (Hendrycks and et al., 2021), DeepSeek-R1 achieves superior accuracy on these complex problems, compared with ChatGPT and Gemini, among others.
@@
See more comparison in DeepSeek-AI (2024, 2025); Zuo and et al. (2025); Arrieta and et al. (2025). Of course, a sophisticated LLM has many aspects (e.g., Infrastructure, Architecture, Performances, Costs) and can achieve many tasks. The tasks discussed above are only a small part of what an LLM can deliver. It is desirable to have a more comprehensive and in-depth comparison. Seemingly, such a comparison may take a lot of time and efforts, but some interesting discussions have already appeared on the internet and social media (e.g., Ramadhan (2025)).
We are especially interested in how accurate an LLM is in prediction. Despite that there are a long list of literature on this topic (see, for example, Friedman et al. (2001)), using LLM for prediction still has advantages: while a classical approach may need a reasonable large training sample, an LLM can work with only a prompt. Along this line, a problem of major interest is how DS compares to other LLMs in terms of prediction accuracy. In this paper, we consider the two classification settings as follows.
  • Authorship classification (AC). Determine whether a document is human generated (hum), or AI-generated (AI), or human-generated but with AI-editing (humAI).
  • Citation classification (CC). Given an (academic) citation and the small piece of text surrounding it, determine which type the citation is (see below for the $4$ types of citations).
@@ 1.1 Authorship classification
In the past two years, AI-generated text content started to spread quickly, influencing the internet, workplace, and daily life. This raises a problem: how to differentiate AI-authored content from human-authored content (Kreps et al., 2022; Danilevsky et al., 2020).
The problem is interesting for at least two reasons. First, the AI-generated content may contain harmful misinformation in areas such as health care, news, and finance (Kreps et al., 2022), and the spread of fake and misleading information may threaten the integrity of online resources. Second, understanding the main differences between human-generated and AI-written content can significantly help improve AI language models (Danilevsky et al., 2020).
@@
For experiments, we propose to use the recent MADStat data set (Ji et al., 2022; Ke et al., 2024). MADStat is a large-scale data set on statistical publications, consisting of the bibtex and citation information of 83,331 papers published in 36 journal in statistics and related field, spanning 1975-2015. The data set is available for free download (see Section 2 for the download links).
We propose a general recipe where we use the LLMs and MADStat to generate new data sets for our study. We start by selecting a few authors and collecting all papers authored by thesm in the MADStat. For each paper, the MADStat contains a title and an abstract.
@@
Seemingly, using this recipe, we can generate many different data sets. These data sets provide a useful platform where we can compare different classification approaches, especially the $5$ LLMs.
Remark 1. (The MadStatAI data set). In Section 2.2, we fix $15$ authors in the MADStat data set (see Table 2 for the list) and generate a data set containing 582 abstract triplets (each triplet contains three abstracts: hum, AI, and humAI) following the recipe above. For simplicity, we call this data set the MadStatAI.
Once the data set is ready, we apply each of the $5$ LLMs above for classification, with the same prompt. See Section 2.1 for details. Note that aside from LLMs, we may apply other algorithms to this problem (Solaiman et al., 2019; Zellers et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2020; Fagni et al., 2021; Adelani et al., 2020; Kashtan and Kipnis, 2024). However, as our focus in this paper is to compare DeepSeek with other LLMs, we only consider the $5$ LLM classifiers mentioned above.
@@ 1.2 Citation classification
When a paper is being cited, the citation could be significant or insignificant. Therefore, to evaluate the impact of a paper, we are interested in not only how many times it is being cited, but also how many significant citations it has. The challenge is, while it is comparably easier to count the raw citations of a paper (i.e., by Google Scholar, Web of Science), it is unclear how to count the number of ‘significant’ citations of a paper.
@@
Here, we have two challenges. First, it is unclear how many different types academic citations may have and what these types are. Second, we do not have a ready-to-use data set.
@@
  • “Fundamental ideas (FI)”
  • “Technical basis (TB)”,
  • “Background (BG)”,
  • “Comparison (CP)”.
Below for simplicity, we encode the four types as “1”, “2”, “3”, “4”. Note that the first two types are considered as significant, and the other two are considered as comparably less significant. See Section 2.2 for details.
Second, with substantial efforts, we have collected from scratch a new data set by ourselves, which we call the CitaStat. In this data set, we downloaded all papers in $4$ representative journals in statistics between 1996 and 2020 in PDF format. These papers contain about $360K$ citation instances. For our study, we selected $n=3000$ citation instances. For each citation,
@@ 1.3 Results and contributions
  • In terms of cost, Claude is much more expensive for a customer than other approaches. For example, for CC1 and CC2 altogether, Claude costs \$$\$12.30$, Llama costs \$$\$1.2$, and the other three methods (DeepSeek, Gemini and GPT) cost no more than \$$\$0.3$.
@@
Table 1: The rankings of all $5$ LLM approaches in terms of error rates.
@@ 2 Main results
In this section, we describe our numerical results on the two problems, authorship classification and citation classification, and report the performances of all 5 LLMs.
@@ 2.1 Authorship classification
The MADStat (Ji et al., 2022; Ke et al., 2024) contains over 83K abstracts, but it is time-consuming to process all of them. [^2] MADStat stands for Multi-Attributed Dataset on Statisticians. MADStat is available for free download at http://zke.fas.harvard.edu/MADStat.html or in the supplementary material of Ji et al. (2022). We selected a small subset as follows: First, we restricted to authors who had over 30 papers in MADStat. Second, we randomly drew 15 authors from this pool by sampling without replacement. Each time a new author was selected, we checked if he/she had co-authored papers with previously drawn authors; if so, we deleted this author and drew a new one, until the total number of authors reached 15. Finally, we collected all 15 authors’ abstracts in MADStat. This gave rise to a data set with 582 abstracts in total (see Table 2).
@@
Both variants are authored by AI, but they look differently. The AI version is usually significantly different from the original abstract, so the ‘human versus AI’ classification problem is comparably easier. For example, the left panel of Figure 1 is a comparison of the length of the original abstract with that of its AI version. The length of human-written abstracts varies widely, while the length of AI-generated ones is mostly in the range of 100-200 words. The humAI version is much closer to the original abstract, typically only having local word replacements and mild sentence re-structuring. In particular, its length is highly correlated with the original length, which can be seen in the right panel of Figure 1.
Figure 1: Comparison of the lengths of human-generated and AI-generated abstracts
@@
Note that comparing with classification approaches (e.g., SVM, Random Forest (Friedman et al., 2001)), an advantage of using an LLM for classification is that, we do not need to provide any training sample. All we need is to input the LLM with a prompt.
@@
We also investigate the similarity of predictions made by different LLMs. For each pair of LLMs, we calculate the percent of agreement on predicted labels, in both the ‘human versus AI’ (AC1) setting and ‘human versus humAI’ (AC2) settings. The results are in Figure 3. For both settings, Gemini-1.5-flash, GPT-4o-mini, and Llama-3.1-8b have extremely high agreements with each other. This is because all three models predict ‘human-written’ for the majority of samples. DeepSeek-R1 and Claude are different from the other three, and they have 64% and 70% agreements with each other in two settings, respectively.
Figure 3: The prediction agreement among 5 LLMs in detecting AI from human texts. Left: ‘human versus AI’ (AC1). Right: ‘human versus humAI’ (AC2). Take the cell on the first row and second column (left panel) for example: for $70\%$ of the samples, the predicted outcomes by DeepSeek-R1 and Claude-3.5-sonnet are exactly the same.
@@ 2.2 Citation classification
The MADStat only contains meta-information and abstracts, rather than full papers. We created a new data set, CitaStat, by downloading full papers and extracting the text surrounding citations. Specifically, we restricted to those papers published during 1996-2020 in 4 representative journals in statistics: Annals of Statistics, Biometrika, Journal of the American Statistical Association, and Journal of the Royal Statistical Society Series B. We wrote code to acquire PDFs from journal websites, convert PDFs to plain text files, and then extracted the sentence (using SpaCy, a well-known python package for sentence parsing) containing each citation (we call it a ‘citation instance’). There are over 367K citation instances in total. We randomly selected $n=3000$ of them and manually labeled them using one of the following four categories:
@@
These definitions were inspired by existing studies on citation types (Moravcsik and Murugesan, 1975; Teufel et al., 2006; Dong and Schäfer, 2011). Occasionally, two categories may overlap. For example, a reference is cited as providing a fundamental idea, and a comparison with it is also stated in the same sentence. In this case, we label it as ‘Fundamental Idea (FI)’, to highlight that this is more important than a general comparison. There were 20 citation instances for which manual labeling gave ‘not sure’. We removed them and obtained $n=2980$ labeled samples (see Table 4).
@@
  • (CC2). The 2-class classification problem: We re-combine the four categories into two, where ‘Fundamental Idea’ and ‘Technical Basis’ are considered as ‘Significant (S)’, and Background and Comparison are considered as ‘Incidental (I)’. [^3] As mentioned, when a citation potentially belongs to two categories (e.g., ‘Fundamental Idea’ and ‘Comparison’), we always manually label it to the more ‘significant’ one (e.g., ‘Fundamental Idea’). This prevents mis-interpreting important comparisons as ‘incidental’ citations in the manual labeling. Given the textual content of a citation, we aim to predict whether it is a ‘Significant (S)’ citation.
For each of the 5 LLMs, we used prompts to get classification decisions. Unlike the previous authorship classification problem, the class definitions in this problem are not common knowledge and need to be included in the prompt. In the 2-class problem, we use the prompt as in Figure 4. It provides definitions, examples, and how the four classes are re-combined into two, aiming to convey to the LLM as much information as possible. The prompt for the 4-class problem is similar, except that the description of grouping 4 classes into 2 is removed and the requirement for output format is modified (see Figure 4).
@@
We examined the performances of all 5 LLMs. Since the runtime of DeepSeek-R1 is much longer than other methods, we only implemented it on 149 randomly selected samples (which included 5% of all samples and maintained the same class fractions as the full data set) to evaluate its classification error rate. Meanwhile, we ran DeepSeek-V3, an earlier-released version of DeepSeek-R1, in all samples. The results are reported in Table 5.
For the 4-class problem, Claude-3.5-sonnet has the lowest error rate at 0.327, followed closely by Gemini-1.5-flash at 0.347. DeepSeek-R1 outperforms DeepSeek-V3 but underperforms other methods except Llama-3.1-8b. For the 2-class problem, Claude-3.5-sonnet still achieves the best performance with an error rate of 0.261. DeepSeek-R1 is the second best, with an error rate of 0.275. Gemini-1.5-flash is worse than DeepSeek-R1 but slightly better than DeepSeek-V3. In summary, Claude-3.5-sonnet is the winner in terms of error rate.
Regarding the runtime, GPT-4o-mini and Gemini-1.5-flash are the fastest (especially, GPT-4o-mini only spent 15 minutes). DeepSeek-V3 and Llama-3.1-8b are relatively slow, requiring a few hours. Regarding the cost, DeepSeek-V3 is the cheapest one, and Claude-3.5-sonnet is significantly most expensive than the other methods.
@@
Claude-3.5-sonnet 0.327 0.261 1-2 h \$ 12.30 USD
DeepSeek-V3 0.432 0.332 3-4 h ¥ 0.60 RMB
[…]
Gemini-1.5-flash 0.347 0.313 25 min \$ 0.12 USD
GPT-4o-mini 0.363 0.371 15 min \$ 0.30 USD
Llama-3.1-8b 0.576 0.457 4-5 h \$ 1.20 USD
[…]
Additionally, in the 4-class setting, we divided all samples into three groups according to the average prediction error by 5 LLMs (this is an average of 5 binary values; we exclude DeepSeek-R1 here, as we don’t have the results on all samples). The lowest 30%, middle 40%, and highest 30% are called the Easy, Medium, and Difficult case, respectively. Table 6 shows the per-group error rates of all 5 LLMs. In the Easy case, the error rates of all methods except Llama-3.1-8b are less than 0.01. In the Difficult case, all methods perform poorly, with GPT-4o-mini attaining the lowest error rate at 0.832 and DeepSeek-V3 attaining the highest at 0.956. In the Medium case, Claude-3.5-sonnet performs well with an error rate of 0.177, followed by Gemini-1.5-flash with a similar error rate 0.211. Llama-3.1-8b shows a notably higher error rate of 0.732.
Claude-3.5-sonnet DeepSeek-V3 Gemini-1.5-flash GPT-4o-mini Llama-3.1-8b
Easy .001 .007 .004 .010 .063
Medium .177 .358 .211 .277 .732
Difficult .851 .956 .870 .832 .881
@@
Table 6: The per-group error rates for 5 LLMs in citation classification.
@@
Finally, we investigate the similarity of predictions made by different LLMs. For DeepSeek-R1, since we only implemented it on 5% of samples, we excluded it in the comparison. For each pair of the remaining 5 LLMs, we calculated the percent of agreement on predicted labels, in both the 4-class and 2-class settings. The results are given in Figure 5.
@@
In the 4-class citation classification setting (CC1), except for Llama-3.1-8b, the percent of agreement between any other two LLMs is above $70\%$. Especially, Claude-3.5-sonnet and Gemini-1.5-flash have the highest level of agreement at $77\%$. Llama-3.1-8b has relatively low agreements with other models. In the 2-class setting, there is a high level of agreement among the 3 LLMs: Claude-3.5-sonnet, DeepSeek-V3, and Gemini-1.5-flash; especially, the agreement between the last two is 83%. Still, Llama-3.1-8b has relatively low agreements with other models. In summary, Gemini-1.5-flash and Claude-3.5-sonnet have a consistently high agreement. For DeepSeek-V3, it has the highest agreement with Gemini-1.5-flash, and its level of agreement with GPT-4o-mini is only in the medium rage.
Figure 5: The prediction agreement among 5 LLMs in citation classification.

### v3

@@ 1 Introduction
In early 2025, DeepSeek (DS), a recent large language model (LLM) (DeepSeek; deepseekai2025deepseekr1incentivizingreasoningcapability), has shaken up the AI industry. Since its latest version was released on January 20, 2025, DS has made the headlines of news and social media, shot to the top of Apple Store’s downloads, stunning investors and sinking some tech stocks including Nvidia.
@@
  • In DS-math, the author showed that over $30$ challenging mathematical problems derived from the MATH dataset (MATHdata), DeepSeek-R1 achieves superior accuracy on these complex problems, compared with ChatGPT and Gemini, among others.
@@
See more comparison in DeepSeek; deepseekai2025deepseekr1incentivizingreasoningcapability; DeepSeek1; DeepSeek2. Of course, a sophisticated LLM has many aspects (e.g., Infrastructure, Architecture, Performances, Costs) and can achieve many tasks. The tasks discussed above are only a small part of what an LLM can deliver. It is desirable to have a more comprehensive and in-depth comparison. Seemingly, such a comparison may take a lot of time and efforts, but some interesting discussions have already appeared on the internet and social media (e.g., DS-comparison).
We are especially interested in how accurate an LLM is in prediction. Despite a long list of statistical and machine learning algorithms for prediction (ESL), using LLM for prediction still has advantages: A classical approach may need a reasonably large training sample, but an LLM can work with only a prompt. Along this line, a problem of major interest is how DS compares to other LLMs in terms of prediction accuracy. In this paper, we consider two classification settings as follows.
  • Authorship classification (AC). Determine whether a document is human-generated (hum), or AI-generated (AI), or human-generated but with AI-editing (humAI).
  • Citation classification (CC). Given an academic citation and the surrounding context, determine which type the citation is (see below for the $4$ pre-defined types).
There are several reasons why we select these two problems for our study. First, these problems are of broad interests. Detection of AI-generated content is an emerging hot topic (liu2019roberta; ippolito2020automatic; solaiman2019release; mitchell2023detectgpt). It enables the AI users to be aware of potential misinformation and incorrect advice in AI-generated content, especially for sensitive areas such as healthcare (brown2020language). Citation analysis is crucial for measuring researchers’ scientific impact (hirsch2005index). For more accurate citation analysis, it is desirable to classify citations into refined categories (teufel2006automatic; li2013towards; dong2011ensemble; pride2020authoritative). Second, these tasks do not require any specific knowledge, making it a fair comparison for LLMs trained for general purposes. Last but not the least, the two problems have the desirable difficulty level for comparing different LLMs: they are neither too easy nor too hard, with the error rates of different LLMs spanning a reasonably wide range.
@@ 1.1 Authorship classification
In the past two years, AI-generated text content started to spread quickly, influencing the internet, workplace, and daily life. This raises a problem: how to differentiate AI-authored content from human-authored content (brown2020language; danilevsky2020survey).
This problem is important for at least two reasons. First, AI-generated content can contain harmful misinformation in critical domains such as healthcare, news, and finance (brown2020language), potentially threatening the integrity of online resources. Second, identifying key differences between human- and AI-generated content can provide valuable insights for improving AI language models (danilevsky2020survey).
@@
For experiments, we propose to use the recent MADStat data set (JBES; TextReview). MADStat is a large-scale data set on statistical publications, consisting of the bibtex and citation information of 83,331 papers published in 36 journal in statistics and related field, spanning 1975-2015. The data set is available for free download (see Section 2 for the download links).
We propose a general recipe where we use the LLMs and MADStat to generate new data sets for our study. We start by selecting a few authors and collecting all papers authored by them in the MADStat. For each paper, the MADStat contains a title and an abstract.
@@
Using this recipe, we can generate many different data sets. These data sets provide a useful platform for us to compare different classification approaches, especially the $5$ LLMs.
Remark 1 (The MadStatAI data set). In Section 2.2, we fix $15$ authors in the MADStat data set (see Table 2 for the list) and generate a data set containing 582 abstract triplets (each triplet contains three abstracts: hum, AI, and humAI) following the recipe above. For simplicity, we call this data set the MadStatAI.
Once the data set is ready, we apply each of the $5$ LLMs above for classification, with the same prompt. See Section 2.1 for details. Note that aside from LLMs, we may apply other algorithms to this problem (solaiman2019release; zellers2019defending; gehrmann2019gltr; ippolito2020automatic; fagni2021tweepfake; adelani2020generating; alonAI; mitchell2023detectgpt). However, since our focus in this paper is to compare DeepSeek with other LLMs, we only consider the $5$ LLM classifiers mentioned above.
@@ 1.2 Citation classification
When a paper is being cited, the citation could be significant or insignificant. Therefore, to evaluate this paper’s impact, we are interested in knowing not only how many times it has been cited, but also how many of those citations are significant. While counting the total number of citations is relatively straightforward using tools like Google Scholar or Web of Science, identifying and quantifying ‘significant’ citations remains a challenging task.
@@
We face two key challenges here. First, the types of different academic citations remain unclear. Second, there is no readily available dataset for analysis.
@@
[…]
For convenience, we encode the four types as “1”, “2”, “3”, “4”. The first two types are considered as significant, and the other two are considered as comparably less significant. See Section 2.2 for details.
Second, with substantial efforts, we have collected from scratch a new data set by ourselves, which we call the CitaStat. We downloaded all papers in $4$ representative journals in statistics between 1996 and 2020 in PDF format. These papers contain about $360,000$ citation instances. For our study, we selected $n=3000$ citation instances. For each citation:
@@ 1.3 Results and contributions
  • In terms of cost, Claude is much more expensive for a customer than other approaches. For example, for CC1 and CC2 altogether, Claude costs $\mathdollar 12.30$, Llama costs $\mathdollar 1.2$, and the other three methods (DeepSeek, Gemini and GPT) cost no more than $\mathdollar 0.3$.
@@
Table 1: The rankings of all 55 LLM approaches in terms of error rates.
@@
Additionally, we have investigated the temporal stability of LLMs, the comparison with traditional classification methods, and the use of different LLMs for generating AI-content.
  • In terms of short-term and long-term variations across multiple runs, Claude, DeepSeek, and Gemini are stable (with more than 90% of agreement on the predicted labels between two separate runs, in the CC1 and CC2 experiments), GPT is moderately stable (with 80-90% agreement), and Llama is unstable (with around 50% agreement).
  • A traditional classification method, the Higher Criticism classifier (DJ08) significantly outperforms all the LLM approaches. Inspired by this observation, we propose a hybrid classifier that combines the strengths of HC and an LLM. In the AC1 experiment, the hybrid of HC and Claude achieves the lowest error rate among all classifiers considered in this paper.
  • Note that we use one LLM to generate the AI content and another LLM to perform classification, and an interesting question is whether it is critical to use the same LLM in the two components. For classification without feature selection by HC, this is not critical. However, for classification with feature selection by HC, we do observe a much better result when we use the same LLM in the two components.
@@ 2 Main results
In Sections 2.1-2.2, we describe our experiments on authorship classification and citation classification, and report the results. In Section 2.3, we investigate the temporal stability of LLMs. In Section 2.4, we compare the performances of LLMs with those of statistical classification approaches, and propose a way for combining the strengths of both.
@@ 2.1 Authorship classification
The MADStat (JBES; TextReview) contains over 83K abstracts, but it is time-consuming to process all of them. [^1] MADStat stands for Multi-Attributed Dataset on Statisticians. MADStat is available for free download at http://tracyke.net/MADStat.html or in the supplementary material of JBES. We selected a small subset as follows: First, we restricted to authors who had over 30 papers in MADStat. Second, we randomly drew 15 authors from this pool by sampling without replacement. Each time a new author was selected, we checked if he/she had co-authored papers with previously drawn authors; if so, we deleted this author and drew a new one, until the total number of authors reached 15. Finally, we collected all 15 authors’ abstracts in MADStat. This gave rise to a data set with 582 abstracts in total (see Table 2).
@@
Both variants are authored by AI, but they differ in appearance. The AI version is often significantly different from the original abstract, so the ‘human versus AI’ classification problem is easier. For example, the left panel of Figure 1 is a comparison of the length of the original abstract with that of its AI version. The length of human-written abstracts varies widely, while the length of AI-generated ones is mostly in the range of 100-200 words. The humAI version is much closer to the original abstract, typically only having local word replacements and mild sentence re-structuring. In particular, its length is highly correlated with the original length, which can be seen in the right panel of Figure 1.
Figure 1: Comparison of the lengths of human-generated and AI-generated abstracts. The x-axis is the length of an original abstract, and the y-axis is the length of its AI counterpart (left panel) or humAI counterpart (right panel).
@@
Note that comparing with classification approaches (e.g., SVM, Random Forest (ESL)), an advantage of using an LLM for classification is that, we do not need to provide any training sample. All we need is to input the LLM with a prompt.
@@
Remark 2: The relatively long runtime of DeepSeek-R1 is primarily due to its design. Compared to DeepSeek-V3 (the version of DeepSeek without reasoning) and other models, DeepSeek-R1 incorporates more complex reasoning chains, multi-step verification processes, and additional internal deliberation layers. While these design choices enhance the model’s accuracy, they also introduce significant computational overhead, making DeepSeek-R1 less practical for time-sensitive or high-throughput applications.
@@
We also investigate the similarity of predictions made by different LLMs. For each pair of LLMs, we calculate the percent of agreement on predicted labels, in both the ‘human versus AI’ (AC1) setting and ‘human versus humAI’ (AC2) settings. The results are in Figure 3. For both settings, Gemini-1.5-flash, GPT-4o-mini, and Llama-3.1-8b have extremely high agreements with each other. This is because all three models predict ‘human-written’ for the majority of samples. DeepSeek-R1 and Claude-3.5-sonnet are different from the other three, and they have 64% and 70% agreements with each other in two settings, respectively.
Figure 3: The prediction agreement among 5 LLMs in detecting AI from human texts. Left: ‘human versus AI’ (AC1). Right: ‘human versus humAI’ (AC2). Take the cell on the first row and second column (left panel) for example: for $64\%$ of the samples, the predicted outcomes by Claude-3.5-sonnet and DeepSeek-R1 are exactly the same.
Remark 3. We briefly explain why an LLM succeeds (or fails) on the above task. Existing studies have shown that AI-generated text and human-written text have different characteristics (e.g., reinhart2025llms). For instance, an LLM may use certain words or sentence structures more frequently than human. Such differences can be detected by analyses of word frequencies or sentence-level likelihoods. In our experiments, especially in AC2, we find that word frequencies contain important signals for distinguishing the AI-generated and human-generated texts (see also Section 2.4). In light of such an observation, we conjecture that an LLM performs well on our tasks for its internal layers activated by the prompt capture some important information of the word frequencies.
Remark 4 (Generation of AI content): To obtain AI-generated content, a popular approach is to pick an LLM, feed it with a small number of words of a document (called the initial words), and then ask it to produce a complete document (jawahar2020automatic). The approach we use GPT-4o-mini to generate paper abstracts follows a similar spirit, where we treat the paper title as the “initial words.” Alternatively, we may feed the whole paper to the LLM and ask it to generate the abstract. As the LLM gets to see more information, this approach may produce an abstract closer to the human-written one. However, such an approach is practically infeasible: Most papers in MADStat are long, and it may exceed the token limit of the LLM even if we only include the introduction of the paper. For this reason, we stick to our current approach above. While our focus is on abstracts generated by GPT-4o-mini, we also investigated abstracts generated by other LLMs; see Section 2.5.
@@ 2.2 Citation classification
The MADStat only contains meta-information and abstracts, rather than full papers. We created a new dataset, CitaStat, by downloading full papers and extracting the text surrounding citations. Specifically, we restricted to those papers published during 1996-2020 in 4 representative journals in statistics: Annals of Statistics, Biometrika, Journal of the American Statistical Association, and Journal of the Royal Statistical Society Series B. We wrote code to acquire PDFs from journal websites, convert PDFs to plain text files, and extract the sentence (using SpaCy, a well-known python package for sentence parsing) containing each citation (we call it a ‘citation instance’). There are over 367K citation instances in total. We randomly selected $n=3000$ of them and manually labeled them using one of the following four categories:
@@
These definitions were inspired by existing studies on citation types (moravcsik1975some; teufel2006automatic; dong2011ensemble). Occasionally, two categories may overlap. For example, a reference is cited as providing a fundamental idea, and a comparison with it is also stated in the same sentence. In this case, we label it as ‘Fundamental Idea (FI)’, to highlight that this is more important than a general comparison. In addition, TB and FI are considered important citations—either the citing paper directly uses the method from the cited work (TB), or the cited work inspires the core contribution of the citing paper (FI). In contrast, CP and BG do not require the cited work to play a critical role in the citing paper. For example, in the sentence “Another way of determining the number of neuron pairs is to follow Medeiros and Veiga (2000b) and Medeiros et al. (2002)”, there is no indication that the methodology in the citing paper is inspired by the cited work and the cited work is only mentioned as alternative approaches to compare. This aligns with our definition of CP, and thus we label it as CP rather than TB.
There were 20 citation instances for which manual labeling gave ‘not sure’. We removed them and obtained $n=2980$ labeled samples (see Table 4).
@@
  • (CC2). The 2-class classification problem: We re-combine the four classes into two, where ‘Fundamental Idea’ and ‘Technical Basis’ are considered as ‘Significant (S)’, and ‘Background’ and ‘Comparison’ are considered as ‘Incidental (I)’. [^2] As mentioned, when a citation potentially belongs to two categories (e.g., ‘Fundamental Idea’ and ‘Comparison’), we always manually label it to the more ‘significant’ one (e.g., ‘Fundamental Idea’). This prevents mis-interpreting important comparisons as ‘incidental’ citations in the manual labeling. Given the textual content of a citation, we aim to predict whether it is a ‘Significant (S)’ citation.
We used prompts to get classification decisions. Unlike the previous authorship classification problem, the class definitions in this problem are not common knowledge and need to be included in the prompt. In the 2-class problem, we use the prompt as in Figure 4. It provides definitions, examples, and how the four classes are re-combined into two, aiming to convey to the LLM as much information as possible. The prompt for the 4-class problem is similar, except that the description of grouping 4 classes into 2 is removed and the requirement for output format is modified (see details in the caption of Figure 4).
@@
We examined the performances of all 5 LLMs. Since the runtime of DeepSeek-R1 is much longer than other methods, we also ran DeepSeek-V3, an earlier-released version of DeepSeek-R1, to evaluate its classification error rate. The results are reported in Table 5.
For the 4-class problem, Claude-3.5-sonnet achieves the lowest error rate at 0.327, followed by Gemini-1.5-flash (0.347) and GPT-4o-mini (0.363). DeepSeek-R1 performs moderately, with an error of 0.380, which is better than DeepSeek-V3 (0.432) and Llama-3.1-8b (0.576) but inferior to the other three models. For the 2-class problem, Claude-3.5-sonnet still has the best error rate (0.261). DeepSeek-R1 is the second best (0.291), outperforming Gemini-1.5-flash (0.313), DeepSeek-V3 (0.332), and GPT-4o-mini (0.371). In summary, Claude-3.5-sonnet achieves the best performance on both problems. DeepSeek-R1, although not the best, shows notable improvement over DeepSeek-V3.
Regarding the runtime, GPT-4o-mini and Gemini-1.5-flash are the fastest, taking only 15 and 25 minutes respectively. DeepSeek-V3 and Llama-3.1-8b require several hours. DeepSeek-R1 is the slowest, taking 60-65 hours (reasons of the slow runtime of DeepSeek-R1 have been discussed in Remark 2). Regarding the cost, DeepSeek-V3 is the cheapest, and Claude-3.5-sonnet is significantly more expensive than the other methods.
@@
Claude-3.5-sonnet 0.327 0.261 1-2 h 12.30 USD
DeepSeek-V3 0.432 0.332 3-4 h 0.60 RMB
DeepSeek-R1 0.380 0.291 60-65 h 11.45 RMB
Gemini-1.5-flash 0.347 0.313 25 min 0.12 USD
GPT-4o-mini 0.363 0.371 15 min 0.30 USD
Llama-3.1-8b 0.576 0.457 4-5 h 1.20 USD
@@
Table 5: The error rate, runtime and cost of 6 LLMs for citation classification.
Additionally, in the 4-class setting, we divided all samples into three groups according to the average prediction error by 6 LLMs (i.e., average of 6 binary values). The lowest 30%, middle 40%, and highest 30% are called the Easy, Medium, and Difficult case, respectively. Table 6 shows the per-group error rates of all 6 LLMs. In the Easy case, the error rates of all methods except Llama-3.1-8b are less than 0.01. In the Difficult case, all methods perform poorly, with GPT-4o-mini attaining the smallest error rate at 0.829 and DeepSeek-V3 attaining the biggest at 0.951. In the Medium case, Claude-3.5-sonnet performs well with an error rate of 0.181, followed by Gemini-1.5-flash with a similar error rate 0.211. Llama-3.1-8b shows a notably higher error rate of 0.727. Compared to DeepSeek-V3, DeepSeek-R1 performs better in all three cases.
Model Easy Medium Difficult
Claude-3.5-sonnet .006 .181 .841
DeepSeek-V3 .004 .364 .951
DeepSeek-R1 .003 .236 .947
Gemini-1.5-flash .006 .211 .869
GPT-4o-mini .009 .280 .829
Llama-3.1-8b .077 .727 .872
Table 6: The per-group error rates for 6 LLMs in citation classification.
@@
Finally, we investigate the similarity of predictions made by different LLMs. For each pair of the 6 LLMs, we calculated the percent of agreement on predicted labels, for both the 4-class and 2-class settings. The results are given in Figure 5.
@@
Noting that DeepSeek-V3 and DeepSeek-R1 are only different versions of the same LLM, we first exclude DeepSeek-V3 and check the agreement among the remaining 5 LLMs. In the 4-class setting (CC1), except for Llama-3.1-8b, the percent of agreement between any other two models is above $70\%$. Especially, each two of Claude-3.5-sonnet, DeepSeek-R1, and Gemini-1.5-flash have a percent of agreement above $75\%$. Llama-3.1-8b has relatively low agreements with other models. In the 2-class setting (CC2), the results are similar: high agreement among Claude-3.5-sonnet, DeepSeek-R1, and Gemini-1.5-flash, and low agreement between Llama-3.1-8b and other models. A notable difference is that GPT-4o-mini has less agreement with other models in the 2-class setting than in the 4-class setting. Also, DeepSeek-V3 agrees with DeepSeek-R1 with most of the entries in classification (82% in the 4-class setting and 84% in the 2-class setting).
Figure 5: The prediction agreement among 6 LLMs in citation classification. Left: 4-class citation classification (CC1). Right: 2-class citation classification (CC2). Take the cell on the first row and second column (left panel) for example: for $73\%$ of the samples, the predicted outcomes by Claude-3.5-sonnet and DeepSeek-V3 are exactly the same.